3-x/x-5 - 4x-7/5-x=
Answer by nerdybill(7384)

A cellular phone company offers a contract for which the cost C, in dollars, of T minutes of telephoning is given by C=0.25(t-600)+50.95, where it is assumed that t is greater then or equal to 600 minutes. What time keeps the cost between $97.95 and $123.45?
.
When C=97.95
97.95=0.25(t-600)+50.95
solving for t:
97.95=0.25t-150+50.95
97.95=0.25t-99.05
197=0.25t
788 minutes = t
.
When C=123.45
123.45=0.25(t-600)+50.95
123.45=0.25t-150+50.95
123.45=0.25t-99.05
222.5 = 0.25t
890 = t
.
answer: 788 minutes <= t <= 890 minutes